OpenStudy (anonymous):

If f(x) = x + 2, find f(3).

OpenStudy (anonymous):

f(3)=3+2

OpenStudy (anonymous):

This just means that you substitute 3 for x

OpenStudy (anonymous):

:o that was a quick reply and thank you :3

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

Plus 3 for x. f( 3) = ( 3) + 2 f( 3) = 5

OpenStudy (anonymous):

I was late XD